Changes

Jump to navigation Jump to search
4 bytes removed ,  02:01, January 6, 2008
rmv unnecessary link
Line 3: Line 3:  
For example, The probability of a tossed fair coin coming up heads is ½ = 0.5 = 50% because there is one favourable outcome (heads) out of 2 possible outcomes (heads and tails).
 
For example, The probability of a tossed fair coin coming up heads is ½ = 0.5 = 50% because there is one favourable outcome (heads) out of 2 possible outcomes (heads and tails).
   −
With a standard deck of 52 [[playing card]]s, there are 13 cards in each suit: spades, hearts, clubs and diamonds. After you shuffle the cards thoroughly, what is the probability that the top card on the deck is a heart? 13 / 52 = 1 / 4 = 0.25 or 25 percent.
+
With a standard deck of 52 playing cards, there are 13 cards in each suit: spades, hearts, clubs and diamonds. After you shuffle the cards thoroughly, what is the probability that the top card on the deck is a heart? 13 / 52 = 1 / 4 = 0.25 or 25 percent.
     
nsTeam1RO, nsTeam1RW, nsTeam1_talkRO, nsTeam1_talkRW
6,781

edits

Navigation menu